OpenStudy (anonymous):

What is the solution for x in the proportion three over five equals twelve over x? 5 10 20 24 @CrashOnce

OpenStudy (crashonce):

again cross multiply: 20 C

OpenStudy (mikurout):

20

OpenStudy (anonymous):

3/5=12/x Cross multiply x= 60/3= 20